You have found the following historical information for the Daniela C mpany Stock price $46.48 2.47 $61.43 2.53 $65 39 2.7 $63.59 2.69 EPS Eamings are expected to grow at 8 percent for the next year. Required: What is the PE ratio for each year? (Do not round intermodiate calculations. Round your answers to 2 decimal places (e.g. 32.16)-) PE ratio Year 1 Year 2 Year 3 Year 4 What is the average PE ratio? (Do not round intermediate calculations. Round your answer to 2 decimal places (e.g. 32.16)) )Average PE
Answer #1

P/E Ratio = Price per share/Earnings per share

Year 1: P/E Ratio = $46.48/$2.47 = 18.82

Year 2: P/E Ratio = $61.43/$2.53 = 24.28

Year 3: P/E Ratio = $65.39/$2.70 = 24.22

Year 4: P/E Ratio = $63.59/$2.69 = 23.64

Average P/E = (18.82 + 24.28 + 24.22 + 23.64)/4 = 22.74